A retarding force f=-2v is acting on a body of mass 10 gram. Find out time taken for its velocity to reduce to 37% of its initial value if initial velocity is 1m/s
1
Expert's answer
2019-08-12T09:56:35-0400

According to the Newton's second law and according to the task

"m\frac{dv(t)}{dt}=-2v(t)"

Solution of the equation with the initial condition

"v(0)=v_0"

is

"v(t)=v_0\exp \left(-\frac{2t}{m}\right)"

So the time taken for its velocity to reduce to 37% of initial value can be found from the equation

"v(t)=0.37v_0"

The solution is

"t=-\frac{m}{2}\ln(0.37)"

And according to the task

"t\approx5\:\mathrm{ms}"

We can see that the result doesn't depend on initial velocity.